# Vital Insight: Assisting Experts’ Context-Driven Sensemaking of Multi-modal Personal Tracking Data Using Visualization and Human-In-The-Loop LLM

### v3

@@ 4.4. Alignment with interview insights
We designed VI-1 based on insights from the initial interviews. In this section, we briefly describe how the features align with those insights. From the interviews, we identified four key sensemaking scenarios that guided the feature design. To address the scenario of contextualizing data based on personalized schedules and self-reported data, we incorporated user check-in and profile modules to provide direct information about the person. Additionally, important contexts such as location labels are integrated into the visualization to aid experts in contextualizing the data. The LLM also incorporates various contextual information to enhance its inferences. To generating insights on various possible events, the LLM provides daily and hourly summaries to help experts understand events at different levels of granularity. We designed the anomaly panels to specifically address the scenario of capturing and explaining anomalies with a quick glance. Lastly, the question panel offers inspiration for actions and interactions with different stakeholders, providing valuable guidance for experts on specific aspects of the data to inspect further. Additionally, the design of VI-1, incorporating both visualization and LLM inference, aims to assist with both direct representation and indirect inference.
@@ 5.1.1. Data collection
To generate the dashboard used in the user testing session, we used real-world data from a deployment with participants, where we collected: 1) pre-study surveys, 2) passive mobile and wearable sensing data, and 3) voice assistant check-ins. Participants completed surveys on demographics and regular routines before deployment. We provided the participants with a Garmin smartwatch and installed the study app on their phones. Our app directly syncs with the Garmin device and collects their physical activity, location, call logs, app usage, Wi-Fi/Bluetooth connections, IBI, heart rate, accelerometer, step counts and more. Additionally, participants can initiate a chit-chat conversation with an LLM-powered chatbot on the study provided Amazon Alexa to check in on their day. The system has been deployed in prior studies (campbell2023patient, ; mishra2021detecting, ; kunzler2019exploring, ; mishra2022towards, ; aip_2024, ), and runs in the background, continuously storing and uploading data to our secure server. We used a secure, HIPAA-compliant GPT-4 model that does not store data for retraining. Additionally, we ensured that no identifiable data is transmitted to the model. For example, we only send the de-identified location labels such as ‘home’ to the external API services instead of raw GPS location, and manually removed potentially identifiable Wi-Fi names and user check-ins. The dashboard presented to the experts does not contain any identifiable information. We used tokens to provide temporary access to the system during the user testing. These precautions ensure that neither GPT-4 nor the participants can identify the subject of data collection. The individuals who provided data consented to share it with external researchers through an IRB-approved study.
@@ 5.1.2. Study Design
Figure 3. Survey results from the initial user testing (Study 2).
[…]
@@ 5.2.2. General usability and qualitative feedback
Experts loved the visualization aspect of the prototype. For example, P3 specifically mentioned that they appreciated how the steps from both the phone and watch were plotted together. Further, experts found the User Profile module helpful (“The user profile, that’s very powerful(P1)”). Similarly, experts loved the user check-in module and wanted it to be more prominent on the dashboard. Some experts like P1 suggested extracting key information from the user check-in panel and displaying it on the visualization according to the time. For example, P1 suggested it would be helpful “if the dashboard said, Oh, this is when the little boy sat. This is when the brunch happened.” as an addition to the original paragraph. Additionally, P4 and others suggested displaying timestamps along with some semantics for the conversation with chatbot in the user check-in module, other than only as an event on the visualization.
@@
During the user testing sessions, many experts appreciated the rich details in both the visualizations and LLM-generated summaries. Many experts also liked the bullet points in the summaries, noting that they provided a good overview. However, experts suggested that it could be even more effective if there were an additional high-level summary beyond the bullet points, such as “more highlights of the day” (P1). This may also be due to the limited time given for experts to complete the tasks, requiring them to explore the system quickly and make assumptions. This feedback indicates that even experts who focus on detailed information still value a simple summary, like “this is a normal Sunday,” as a starting point.
@@
Many experts found the LLM-generated summaries to be “good” and “beyond expectations.” Experts generally agreed that the system is evidence-based, which aligns with the survey results. Through interaction with the various components, many experts were able to identify the evidence used by the LLM to make its assumptions. Experts also appreciated the different possibilities of events provided by the LLM, and enjoyed the current ways where evidence is generally separated with inference. However, many expressed a desire for the evidence to be more apparent, such as by highlighting the data used in the visualization or indicating when information is derived from user check-ins. P5 mentioned: “I don’t mind making some guess, but I wanted to know where that info (from LLM) is coming from.”
@@
Another interesting insight emerged when examining the anomaly: the differing definitions of what should be considered an anomaly. Many experts noted that the current anomalies are often more like “standout events” rather than true anomalies or alerts. Almost all experts confirmed that there were hardly any anomalies needing immediate concern in current data, and emphasized they would need to better understand the person’s regular patterns and routines to further classify anomalies. This highlights a potential need for the prototype to distinguish more clearly between special events or highlights and genuine anomalies that experts should be concerned about, which we will discuss further in the next section.
@@ 5.2.3. User interaction: Experts’ sensemaking process
In this section, we further examine experts’ thought processes in sensemaking through their self-descriptions (think-aloud) and recorded screen interactions. Experts employ a variety of approaches for sensemaking, with notable commonalities as well as differences. Most experts started their analysis with the visualization (P4/5/6/8/10/11/12/13), while P9 started with the LLM summary, and P1/2/3/7 began with the user profile. Experts quickly switched to the other to verify details, moving back and forth between the visualization and LLM contents. Some experts composed the summary entirely on their own, whereas others opted to copy and paste the LLM inferences into the summary and edit (P8, P13), such as removing information they deemed trivial or unimportant (e.g., WiFi details, specific phone usage patterns) and adding significant details not mentioned in the LLM-generated summary (e.g., sleep patterns, correlations between activity, and heart rate).
To further scaffold their sensemaking process, we first confirmed that experts’ sensemaking process is inherently iterative, moving from high-level insights to detailed analysis and then back to a broader perspective. Through the iterative process, experts integrate multiple modalities to make assumptions. For closely linked modalities, some processes also served as a form of validation. For instance, P11 initially assumed that a decrease in battery level indicated screen time but rejected this assumption on noticing that the phone was locked during that period. Validation across modalities was frequent and varied, with common comparisons between WiFi, location, heart rate, respiration, steps, and activity. Experts also attempted to integrate and compare user check-ins with other sensing modalities. However, this was more difficult due to the current setup, where user check-ins are fully separate from the main visualization. As a result, they expressed the need for the LLM to extract check-in information and display it on the visualization to facilitate easier comparison between modalities.
@@
Figure 4. Design of the new iteration of Vital Insight (VI-2). Top left: user testing input; bottom left: main visualization and LLM detected occurrences; top right: user profile; middle right: Day in a glance generated by LLM; bottom right: user check-in conversation.
[…]
@@ 5.3. Follow-up evaluation on LLM anomaly detection
Through the interviews, we discovered that the LLM’s anomaly detection on a single modality did not align with the experts’ expectations. As a result, we conducted an exploratory analysis to evaluate LLM anomaly detection. We ran the LLM anomaly detection on heart rate and respiration data for a week (11/18/2024 - 11/24/2024), performing 10 tests for each day using the same prompts. We observed significant instability in the detection on several days (Figure.  5). Ideally, all the different colors of the horizontal lines should align for the same day. In our evaluation, although there is some consistency on several days (e.g., heart rate anomaly on 11/20), many outputs are scattered across the 24-hour period. Additionally, there are two significant long heart rate anomaly durations in the initial LLM outputs on 11/23 and 11/19 that are inaccurate. As a result, the LLM anomaly detection results not only misalign with the experts’ mental model but are also highly inconsistent and unstable. This highlights the need for a more expert-instructed anomaly detection process to ensure greater stability.
Figure 5. Time ranges of LLM detected heart rate and respiration anomalies from 11/18/2024 to 11/24/2024. The horizonal lines represents the start and end time of the anomalies, and the different colors represent 10 LLM test runs.
[…]
@@ 6. Final Prototype: Seamless Integration of AI Assistance with Visualization
Based on feedback from Study 2, a major area for improvement was the seamless integration of AI augmentation with the visualization. For instance, experts expect not just a text-based summary, but the ability to extract important information and display it directly within the visualization to fill in the gaps. Regarding trustworthiness, we found that experts struggled to trust the anomaly detection results, especially when they were purely generated by the LLM. As a result, we modified the design to incorporate a more human-in-the-loop process to detect and explain important events, which we defined as “occurrences” in our final prototype. Below, we detail the changes we made to design a final prototype, Vital Insight 2 (VI-2) (Figure.  4).
@@
In addition to the backend improvements on how to identify and explain occurrences, VI-2 introduces a new way of displaying these occurences to experts. Instead of being a separate module (like the anomalies module in VI-1), Occurrences and their LLM-generated explanations are now directly plotted on top of the related raw data streams, represented by orange dots with name of that occurrence. When a user hovers over an orange dot, they will see the title of the occurence, a detailed inference generated by the LLM, and a description of the data sources used by the LLM to generate that inference. This method represents an approach where, instead of providing a stand-alone result like a paragraph in Study 2, we aim to create AI assistance that is seamlessly integrated into the experts’ sensemaking process. The new LLM inferences support the non-chronological sensemaking process of experts as well as provide a source of input for AI assistance to build trust. We envision that this newly updated prototype can better support the collaborative process between experts, data, and AI assistance. We have made the system open-source and publicly accessible on GitHub at [anonymized link].
@@ 7.1.1. Data collection
We used the same sensing system to collect data. Additionally, to determine whether the sensemaking results from researchers were accurate, participants self-reported fine-grained activities through ecological momentary assessments (EMA) multiple times a day, in addition to the daily check-ins with Alexa. The EMA results typically consisted of a few words describing the current activity, such as “going out for lunch.” We used EMAs solely as the ground truth for momentary user behavior to determine the accuracy of the if the sensemaking results from researchers. We did not provide the EMAs to the LLMs or the experts.
@@ 7.1.2. Study Design
We conducted an A/B test to investigate whether LLM assistance in VI-2 is helpful for researchers in performing sensemaking tasks. Condition A used VI-2 with full LLM assistance, while Condition B used VI-2 without LLM assistance. In Condition B, the experts did not have access to the Day-in-a-Glance and the Occurrences explanation. During the experiment, each participant interacted with data from four different days, with two days for each condition, and answered a series of questions. The order of the tasks (A/B conditions and order of days) was randomized to ensure counterbalance and avoid carry-over effect. For each day, researchers first answered three multiple-choice questions related to the person’s day using all the information provided in the system. For example, a sample question could be “What is the person doing at 6:30 pm?”, with several options like “going out for lunch”, “hanging out with friends”, etc. After answering each question, experts rated their experience by completing the Single Usability Metrics survey (SUM) across the dimensions of ease, time, and satisfaction (sauro2005method, ), along with two Likert-scale questions to rate the AI component (only in Condition A, where LLM is enabled). Next, participants wrote a short summary of the person’s day within 3 minutes and completed the same surveys they filled out after answering the previous selection questions. After each task, participants discussed their thought processes as qualitative feedback for the research team. The questions module was integrated directly into the interface during the user testing  4. After the experiment, the participants filled out two System Usability Scale (SUS) to rate the overall usability of LLM enabled and disabled prototypes (brooke1996sus, ), and answered several questions related to the helpfulness of AI assistance. At the end of the study, we asked participants several questions aimed at identifying potential actions based on the four days of data. For example, we asked what insights could be generated about the person, and inquired about the types of health interventions they might design to improve the person’s physical activity levels. Since the correctness of the answers could not be objectively measured during this user testing, we kept this as an open-ended discussion with the experts to gather qualitative insights.
@@
For Study 3, we followed the same recruiting and compensation process as Study 2. We recruited a total of 20 participants, with some overlap with Study 1 and 2 (Table. 1). We recorded all user testing sessions, and followed an open-coding approach to analyze the transcripts and notes as our qualitative analysis. For the quantitative analysis, we calculated various metrics across the two conditions, including task completion time, correctness in answering the sensemaking questions by comparing with EMA ground truth, user-perceived metrics (satisfaction and ease), and the SUS score for the entire system.
Figure 6. Study design of the A/B user testing (Study 3).
[…]
@@ 7.3.1. Descriptive Results
Figure 7. Descriptive results from the A/B user testing (Study 3).
[…]
@@ 7.3.2. Results of A/B test
We additionally conducted statistical tests to evaluate the differences in completing sensemaking tasks (answering questions and writing summary) with and without LLM. For ordinal variables, such as ease and satisfaction measured on a scale from 1 to 5, we used a Cumulative Logit Mixed-effects Model (CLMM) to examine the effect of LLM on participants’ responses. For numerical variables, such as correctness (scored from 0 to 1) and completion time, we used Linear Mixed-Effects Models (LMER). We included random intercepts for id (participant) and nested random effects for date and question, as each date had different questions, accounting for variability in responses within participants and across different dates and questions. Since each participant was only exposed to one condition for the same date, we fitted two models: one that included an interaction term between condition and date to test whether the effect of condition varied across dates, and a simpler model that included only the main effect of condition. We compared the models to assess whether the inclusion of the interaction term improved the model fit, using the Akaike Information Criterion (AIC) value and selected the model with lower AIC value.
@@
The distribution of SUS scores with LLM followed a normal distribution (mean = 73.42, SD = 17.92), while the distribution without LLM was negatively skewed (mean = 70.53, median = 75, SD = 19.59). The mean SUS scores were considered “Good” and “Acceptable” for the LLM-enabled condition, and “OK” and “Marginal High” for the LLM-disabled condition (bangor2009determining, ; bangor2008empirical, ). Using a Wilcoxon Signed-Rank Test, we found that having AI-assistance led significantly higher SUS scores than without AI-assistance($p<0.05$), with a medium effect (effect size = 0.44).

### v4

@@ 4.4. Alignment with Interview Insights
We designed VI-1 based on insights from the initial interviews. In this section, we briefly describe how the features align with those insights. From the interviews, we identified four key sensemaking scenarios that guided the feature design. To address the scenario of contextualizing data based on personalized schedules and self-reported data, we incorporated user check-in and profile modules to provide direct information about the person. Additionally, important contexts such as location labels are integrated into the visualization to aid experts in contextualizing the data. The LLM also incorporates various contextual information to enhance its inferences. To generate insights on various possible events, the LLM provides daily and hourly summaries to help experts understand events at different levels of granularity. We designed the anomaly panels to specifically address the scenario of capturing and explaining anomalies with a quick glance. Lastly, the question panel offers inspiration for actions and interactions with different stakeholders, providing valuable guidance for experts on specific aspects of the data to inspect further. Additionally, the design of VI-1, incorporating both visualization and LLM inference, aims to assist with both direct representation and indirect inference.
@@ 5.1.1. Data Collection
To generate the dashboard used in the user testing session, we used real-world data from a deployment with participants, where we collected: 1) pre-study surveys, 2) passive mobile and wearable sensing data, and 3) voice assistant check-ins. Participants completed surveys on demographics and regular routines before deployment. We provided the participants with a Garmin smartwatch and installed the study app on their phones. Our app directly syncs with the Garmin device and collects their physical activity, location, call logs, app usage, Wi-Fi/Bluetooth connections, IBI, heart rate, accelerometer, step counts and more. Additionally, participants can initiate a chit-chat conversation with an LLM-powered chatbot on the study provided Amazon Alexa to check in on their day. The system has been deployed in prior studies (campbell2023patient, ; mishra2021detecting, ; kunzler2019exploring, ; mishra2022towards, ; aip_2024, ), and runs in the background, continuously storing and uploading data to our secure server. We used a secure, HIPAA-compliant GPT-4 model that does not store data for retraining. Additionally, we ensured that no identifiable data is transmitted to the model. For example, we only send the de-identified location labels such as ‘home’ to the external API services instead of raw GPS location, and manually remove potentially identifiable Wi-Fi names and user check-ins. The dashboard presented to the experts does not contain any identifiable information. We used tokens to provide temporary access to the system during the user testing. These precautions ensure that neither GPT-4 nor the participants can identify the subject of data collection. The individuals who provided data consented to share it with external researchers through an IRB-approved study.
@@ 5.1.2. Study Design
Figure 3. Survey results from the initial user testing (Study 2).:
@@ 5.2.2. General Usability and Qualitative Feedback
Experts loved the visualization aspect of the prototype. For example, P3 specifically mentioned that they appreciated how the steps from both the phone and watch were plotted together. Further, experts found the User Profile module helpful (“The user profile, that’s very powerful(P1)”). Similarly, experts loved the user check-in module and wanted it to be more prominent on the dashboard. Some experts like P1 suggested extracting key information from the user check-in panel and displaying it on the visualization according to the time. For example, P1 suggested it would be helpful “if the dashboard said, Oh, this is when the little boy sat. This is when the brunch happened.” as an addition to the original paragraph. Additionally, P4 and others suggested displaying timestamps along with some semantics for the conversation with the chatbot in the user check-in module, other than only as an event on the visualization.
@@
During the user testing sessions, many experts appreciated the rich details in both the visualizations and LLM-generated summaries. Many experts also liked the bullet points in the summaries, noting that they provided a good overview. However, experts suggested that it could be even more effective if there were an additional high-level summary beyond the bullet points, such as “more highlights of the day” (P1). This may also be due to the limited time given for experts to complete the tasks, requiring them to explore the system quickly and make assumptions. This feedback indicates that even experts who focus on detailed information still value a simple summary, like “this is a normal Sunday” as a starting point.
@@
Many experts found the LLM-generated summaries to be “good” and “beyond expectations.” Experts generally agreed that the system is evidence-based, which aligns with the survey results. Through interaction with the various components, many experts were able to identify the evidence used by the LLM to make its assumptions. Experts also appreciated the different possibilities of events provided by the LLM, and enjoyed the current ways in which evidence is generally separated from inference. However, many expressed a desire for the evidence to be more apparent, such as by highlighting the data used in the visualization or indicating when information is derived from user check-ins. P5 mentioned: “I don’t mind making some guess, but I wanted to know where that info (from LLM) is coming from.”
@@
Another interesting insight emerged when examining the anomaly: the differing definitions of what should be considered an anomaly. Many experts noted that the current anomalies are often more like “standout events” rather than true anomalies or alerts. Almost all experts confirmed that there were hardly any anomalies needing immediate concern in the current data, and emphasized they would need to better understand the person’s regular patterns and routines to further classify anomalies. This highlights a potential need for the prototype to distinguish more clearly between special events or highlights and genuine anomalies that experts should be concerned about, which we will discuss further in the next section.
@@ 5.2.3. User Interaction: Experts’ Sensemaking Process
In this section, we further examine experts’ thought processes in sensemaking through their self-descriptions (think-aloud) and recorded screen interactions. Experts employ a variety of approaches for sensemaking, with notable commonalities as well as differences. Most experts started their analysis with the visualization (P4/5/6/8/10/11/12/13), while P9 started with the LLM summary, and P1/2/3/7 began with the user profile. Experts quickly switched to the other to verify details, moving back and forth between the visualization and the LLM contents. Some experts composed the summary entirely on their own, whereas others opted to copy and paste the LLM inferences into the summary and edit (P8, P13), such as removing information they deemed trivial or unimportant (e.g., WiFi details, specific phone usage patterns) and adding significant details not mentioned in the LLM-generated summary (e.g., sleep patterns, correlations between activity, and heart rate).
To further scaffold their sensemaking process, we first confirmed that experts’ sensemaking process is inherently iterative, moving from high-level insights to detailed analysis and then back to a broader perspective. Through the iterative process, experts integrate multiple modalities to make assumptions. For closely linked modalities, some processes also served as a form of validation. For instance, P11 initially assumed that a decrease in battery level indicated screen time, but rejected this assumption on noticing that the phone was locked during that period. Validation across modalities was frequent and varied, with common comparisons between WiFi, location, heart rate, respiration, steps, and activity. Experts also attempted to integrate and compare user check-ins with other sensing modalities. However, this was more difficult due to the current setup, where user check-ins are fully separate from the main visualization. As a result, they expressed the need for the LLM to extract check-in information and display it on the visualization to facilitate easier comparison between modalities.
@@
Figure 4. Time ranges of LLM detected heart rate and respiration anomalies from 11/18/2024 to 11/24/2024. The horizontal lines represent the start and end time of the anomalies, and the different colors represent 10 LLM test runs.:
@@ 5.3. Follow-up Evaluation on LLM Anomaly Detection
Through the interviews, we discovered that the LLM’s anomaly detection on a single modality did not align with the experts’ expectations. As a result, we conducted an exploratory analysis to evaluate LLM anomaly detection. We ran the LLM anomaly detection on heart rate and respiration data for a week (11/18/2024 - 11/24/2024), performing 10 tests for each day using the same prompts. We observed significant instability in the detection on several days (Figure.  4). Ideally, all the different colors of the horizontal lines should align for the same day. In our evaluation, although there is some consistency on several days (e.g., heart rate anomaly on 11/20), many outputs are scattered across the 24-hour period. Additionally, there are two significant long heart rate anomaly durations in the initial LLM outputs on 11/23 and 11/19 that are inaccurate. As a result, the LLM anomaly detection results not only misalign with the experts’ mental model but are also highly inconsistent and unstable. This highlights the need for a more expert-instructed anomaly detection process to ensure greater stability.
@@ 6. FINAL PROTOTYPE: SEAMLESS INTEGRATION OF AI ASSISTANCE WITH VISUALIZATION
Based on feedback from Study 2, a major area for improvement was the seamless integration of AI augmentation with the visualization. For instance, experts expect not just a text-based summary, but the ability to extract important information and display it directly within the visualization to fill in the gaps. Regarding trustworthiness, we found that experts struggled to trust the anomaly detection results, especially when they were purely generated by the LLM. As a result, we modified the design to incorporate a more human-in-the-loop process to detect and explain important events, which we defined as “occurrences” in our final prototype. Below, we detail the changes we made to design a final prototype, Vital Insight 2 (VI-2) (Figure.  5).
@@
In addition to the backend improvements on how to identify and explain occurrences, VI-2 introduces a new way of displaying these occurrences to experts. Instead of being a separate module (like the anomalies module in VI-1), Occurrences and their LLM-generated explanations are now directly plotted on top of the related raw data streams, represented by orange dots with the name of that occurrence. When a user hovers over an orange dot, they will see the title of the occurrence, a detailed inference generated by the LLM, and a description of the data sources used by the LLM to generate that inference. This method represents an approach where, instead of providing a stand-alone result like a paragraph in Study 2, we aim to create AI assistance that is seamlessly integrated into the experts’ sensemaking process. The new LLM inferences support the non-chronological sensemaking process of experts as well as provide a source of input for AI assistance to build trust. We envision that this newly updated prototype can better support the collaborative process between experts, data, and AI assistance. We have made the system open-source and publicly accessible on GitHub at https://github.com/UbiWell/vital-insight-public-repo.
Figure 5. Design of the new iteration of Vital Insight (VI-2). Top left: user testing input; bottom left: main visualization and LLM detected occurrences; top right: user profile; middle right: Day in a glance generated by LLM; bottom right: user check-in conversation.:
@@ 7.1.1. Data Collection
We used the same sensing system to collect data. Additionally, to determine whether the sensemaking results from researchers were accurate, participants self-reported fine-grained activities through ecological momentary assessments (EMA) multiple times a day, in addition to the daily check-ins with Alexa. The EMA results typically consisted of a few words describing the current activity, such as “going out for lunch.” We used EMAs solely as the ground truth for momentary user behavior to determine the accuracy of if the sensemaking results from researchers. We did not provide the EMAs to the LLMs or the experts.
@@ 7.1.2. Study Design
We conducted an A/B test to investigate whether LLM assistance in VI-2 is helpful for researchers in performing sensemaking tasks. Condition A used VI-2 with full LLM assistance, while Condition B used VI-2 without LLM assistance. In Condition B, the experts did not have access to the Day-in-a-Glance and the Occurrences explanation. During the experiment, each participant interacted with data from four different days, with two days for each condition, and answered a series of questions. The order of the tasks (A/B conditions and order of days) was randomized to ensure counterbalance and avoid a carry-over effect. For each day, researchers first answered three multiple-choice questions related to the person’s day using all the information provided in the system. For example, a sample question could be “What is the person doing at 6:30 pm?”, with several options like “going out for lunch”, “hanging out with friends”, etc. After answering each question, experts rated their experience by completing the Single Usability Metrics survey (SUM) across the dimensions of ease, time, and satisfaction (sauro2005method, ), along with two Likert-scale questions to rate the AI component (only in Condition A, where LLM is enabled). Next, participants wrote a short summary of the person’s day within 3 minutes and completed the same surveys they filled out after answering the previous selection questions. After each task, participants discussed their thought processes as qualitative feedback for the research team. The questions module was integrated directly into the interface during the user testing  5. After the experiment, the participants filled out two System Usability Scale (SUS) to rate the overall usability of LLM-enabled and disabled prototypes (brooke1996sus, ), and answered several questions related to the helpfulness of AI assistance. At the end of the study, we asked participants several questions aimed at identifying potential actions based on the four days of data. For example, we asked what insights could be generated about the person, and inquired about the types of health interventions they might design to improve the person’s physical activity levels. Since the correctness of the answers could not be objectively measured during this user testing, we kept this as an open-ended discussion with the experts to gather qualitative insights.
@@
For Study 3, we followed the same recruiting and compensation process as Study 2. We recruited a total of 20 participants, with some overlap with Study 1 and 2 (Table. 1). We recorded all user testing sessions and followed an open-coding approach to analyze the transcripts and notes as our qualitative analysis. For the quantitative analysis, we calculated various metrics across the two conditions, including task completion time, correctness in answering the sensemaking questions by comparing with EMA ground truth, user-perceived metrics (satisfaction and ease), and the SUS score for the entire system.
Figure 6. Study design of the A/B user testing (Study 3).:
@@ 7.3.1. Descriptive Results
Figure 7. Descriptive results from the A/B user testing (Study 3).:
@@ 7.3.2. Results of A/B Test
We additionally conducted statistical tests to evaluate the differences in completing sensemaking tasks (answering questions and writing summaries) with and without LLM. For ordinal variables, such as ease and satisfaction measured on a scale from 1 to 5, we used a Cumulative Logit Mixed-effects Model (CLMM) to examine the effect of LLM on participants’ responses. For numerical variables, such as correctness (scored from 0 to 1) and completion time, we used Linear Mixed-Effects Models (LMER). We included random intercepts for id (participant) and nested random effects for date and question, as each date had different questions, accounting for variability in responses within participants and across different dates and questions. Since each participant was only exposed to one condition for the same date, we fitted two models: one that included an interaction term between condition and date to test whether the effect of condition varied across dates, and a simpler model that included only the main effect of condition. We compared the models to assess whether the inclusion of the interaction term improved the model fit, using the Akaike Information Criterion (AIC) value and selected the model with lower AIC value.
@@
The distribution of SUS scores with LLM followed a normal distribution (mean = 73.42, SD = 17.92), while the distribution without LLM was negatively skewed (mean = 70.53, median = 75, SD = 19.59). The mean SUS scores were considered “Good” and “Acceptable” for the LLM-enabled condition, and “OK” and “Marginal High” for the LLM-disabled condition (bangor2009determining, ; bangor2008empirical, ). Using a Wilcoxon Signed-Rank Test, we found that having AI-assistance led to significantly higher SUS scores than without AI-assistance($p<0.05$), with a medium effect (effect size = 0.44).